# Visualization of Electronic Health Record Sequences at Scale

### v1

@@ 4.3.2 Results
Our evaluation shows that all nodes have a similar frequency from the first iteration to the last, regardless of MinSize and chunkSize, meaning that the frequency distribution of nodes converges fast for datasets like ours. As we shuffled the datasets, this result is expected for high values of chunkSize. We then do not report on the variable Bad_Statistics further.
@@
We then look at how fast Hysteresis Sort and RSort stabilize in depth and relative position (Figure 9 and Figure 10). We normalized the iteration ID from 0 to 1 to compare all evaluations with different chunkSize. The distribution of the number of stabilized nodes and depths compared to the iteration ID is not normal. Thus, we cannot rely on parametric statistical tests. Following recommendations of [11, 1, 4], we instead rely on estimation techniques with effect-size and confidence intervals (CIs). We bootstrap all the evaluation results per sorting strategy and per depth, resulting in $\textit{size}(\textit{MinSize})\times\textit{size}(\textit{chunkSize})$ different values per bootstrap. Depth starts from 0, which corresponds to the children of the root, as the root is, in our case, an abstract object that we do not visualize in ParcoursVis.
@@
Figure 9 shows that the prefix tree is stable after the progressive algorithm processed $80\%$ (at the right edge of the 95% CI) of the dataset at the maximum tested $\textit{depth}=4$ using the Hysteresis Sort, compared to approximately 90% using RSort. Our results (Figure 10) show strong evidence (statistically significant differences) that Hysteresis Sort and RSort start to diverge at depth=1, with a large effect size for $\textit{depth}\in\{1,2\}$. Because deeper depths depend on shallower ones, the results of RSort for $\textit{depth}=1$ strongly impact the results of deeper depths. This seems less the case for Hysteresis Sort (Figure 9).
@@
We then looked at the stability per node, categorized by their depths. Figure 10 shows strong evidence that Hysteresis Sort outperforms RSort at stabilizing nodes sooner for all $\textit{depth}\geq 1$. In addition, we looked at the stability of nodes categorized by their frequencies. We computed, per node, the difference in stability of Hysteresis Sort compared to RSort for the evaluation, set with $\textit{MinSize}=0$ and $\textit{chunkSize}=100K$, which are the minimum values we tested. We see that Hysteresis Sort and RSort behave the same for most nodes, which we expected, as most nodes do not have siblings with similar frequencies (see $\textit{iteration}=0$). Second, the sorting strategy mostly impacts nodes with low frequency (lower than 200, i.e., representing less than 1% of our population). Third, while Hysteresis Sort strongly improves the stability of most of the remaining nodes, it can degrade the stability of some small nodes for a few iterations. Overall, we conclude on the effectiveness of Hysteresis Sort compared to RSort.
@@
Finally, we wanted to see if the variables MinSize and chunkSize have an effect on the stability. Figure 11 suggests that MinSize strongly impacts the stability of the positions of the nodes by removing small changes. Because the dataset is shuffled, large nodes tend to stabilize sooner than smaller ones, explaining the results. However, this figure shows that MinSize has a stronger impact on the Hysteresis Sort than on the RSort. We do not find any effect of the chunkSize variable on the stability of nodes.
5 Applications
@@
ParcoursVis is applied to two distinct use cases, enabling participants to visualize real-world large datasets with an unprecedented level of scalability. We provide usability insights, focusing on its progressive aspect when used by experts. Participants were able to perform tasks comparable to those in other ESS systems with no perceived latency.
5.1 Non-cancerous prostate adenoma in France
@@
We provided ParcoursVis to referring doctors to explore the non-cancerous prostate adenoma dataset. We received positive feedback. One referring doctor reported that his wife thought he had a new video game because he was spending hours exploring the data interactively. They were excited to be able to explore such a large dataset to eventually find unexpected cases. One of the referring doctors has collected many insights and surprising findings that need inquiry now. For example, they thought that surgery would cure the patients when, in reality, many patients ($18\%$) need to resume treatment and sometimes undergo another surgery.
5.2 Emergency Pathways in Parisian Hospitals
We also used ParcourVis for visualizing patient pathways in the 16 public Emergency Departments (EDs) in Paris. For over 15 years, EDs have faced a steady increase in patient volume, leading to frequent overcrowding and prolonged waiting times. Analyzing patient care pathways offers the potential to identify specific stages where the process becomes delayed, resulting in extended waiting times or medical accidents. Our project deals with 1.5M patient visits over three years. We wrapped the ParcoursVis front end into a widget for use within the secure Jupyter Lab environment provided by Paris Public Hospitals IT, allowing for easy access by our collaborators.
We interviewed, in an ongoing usability study, two professionals interested in analyzing patients’ pathways: an emergency physician highly motivated to find more effective organizational approaches and a computer scientist aiming at achieving high-fidelity simulation of patients’ flow in EDs. None of our participants mentioned any latency issues, not even performance issues. They were able to focus on their professional questions.
@@ 5.2 Emergency Pathways in Parisian Hospitals
We also used ParcoursVis to answer several questions during the project’s meetings, allowing us to quickly characterize and select important pathways to optimize, such as the most frequent ones, the most time-consuming ones, and the most “undesirable” ones.
6 Discussion
6.1 Performance Analysis
@@
The competitive analysis (Section 4.2) validates the linear complexity of our aggregation algorithm, both in the number of threads and low-level events. We aggregate 10 million synthetic patients ($\approx 180M$ low-level events) on our dedicated computer in less than one second. ParcoursVis is meant to be used on a server shared by multiple users; aggregating sequences in the wild takes longer. Moreover, our evaluation relies on short sequences on which we applied a small number of filters. Longer sequences and more filters cause linear slowdowns. PVA solves these issues with almost no overhead.
@@
We filter small nodes to avoid confetti visualizations, but users can configure it. In our evaluation using $\textit{MinSize}=50$ (i.e., 0.0025% of the two million patients dataset), the view aggregating the dataset stabilizes at around half the computation time for the deepest nodes using Hysteresis Sort, 25% sooner than RSort (Figure 11). The MinSize variable thus can strongly increase the speed at which users can work on their data if they are not interested in small nodes.
6.2 Stability
@@
We evaluated the stability of aggregated trees for one use case with an inertia of $20/1080$. For that configuration, we show Hysteresis Sort is better than RSort. We believe this generalizes to other datasets, although the stabilization times may vary depending on the distribution of events. Our Hysteresis Sort algorithm relies on a modified version of the bubble sort algorithm with a theoretical worst-case quadratic complexity. In practice, our sort time is linear because we use a limited number of types to be able to assign them to color effectively. The scalability of ParcoursVis and ESS, in general, remains limited by the number of event types they can handle [49].
6.3 Evaluation Criteria for Progressive Visualization
Progressive systems are relatively new, and we need more evaluation criteria to assess their usability [37].
@@ 6.3 Evaluation Criteria for Progressive Visualization
We did not consider comparing the usability of our system in its current state to a non-progressive system, given the latency we measured, which makes non-progressive systems completely unusable (measures are in the Supplementary Material). During our expert interviews, participants were unaware that the tool used progressive algorithms and did not notice it; we consider this lack of mention a usability success for progressive systems and algorithms. It is a new evaluation criterion.
@@
As explained in Section 2.4, we rely on stability as a proxy for convergence. For update durations limited to tens of seconds, we believe instability is an acceptable proxy for uncertainty, and our expert users did not express any confusion or trust issues about it.
ParcoursVis could also include visualization elements that indicate the threshold for our hysteresis node ordering, coupled with direct user interaction. With this more advanced feature, conducting studies on trust vs. stability in the visualization could be relevant.
6.4 Improving the Scalability of EHR Visualizations
@@
7 Conclusion
We presented ParcoursVis, a PVA tool we designed to explore patients’ care pathways at scale. We described its use, including the history navigation and comparison that are novel for HRS systems. We described its scalable aggregation algorithm and its stabilizing sorting algorithm. We showed quantitative evaluations showing the aggregation algorithm scales linearly with the number of events and inversely with the number of threads. We also showed that our Hysteresis sort improved the stability of the Icicle tree visualization. ParcoursVis can be used effectively to explore patients’ treatments and conditions aggregated at the scale of the largest countries to improve public health based on data.
For a dataset scaled to China’s population, non-progressive aggregation on our machine would take 27s. In contrast, our PVA architecture ensures updates in under 1s, with most tree nodes stabilizing within 10s, enabling exploration of datasets at the scale of countries.
Each new study requires new features, and we are planning to extend ParcoursVis to make it more generic and applicable to new use cases without programming. We also plan to support datasets updated continuously. That would require certain adjustments to the system, specifically running regular pre-processings of the updated database, e.g., every 15 minutes. ParcoursVis should be able to restart immediately when a new dataset is available, but it should avoid interrupting ongoing explorations. By providing our system as open-source, we want to push this research field to a wider audience, relying on PVA to provide scalability and improve public health.
@@ 9 Histogram age_distribution; // age distribution
Listing 4: Structure of a tree Node
The tree contains nodes (LABEL:lst:node), each node $N_{i}$ representing a pathway $P_{i}$ from the root to $N_{i}$:
@@
  • children: The list of events following $P_{i}$. For each event $E_{j}$, we export its frequency, i.e., probability of $P({E_{j}}|{P_{i}})$.
@@
From this tree, we can generate a new patient in three phases: (1) generating a random sequence of high-level events from the tree, (2) adding realistic attributes to the high-level events, and (3) generating a plausible low-level event sequence from the high-level events. While it is common to rely on statistics and machine learning models [17, 48, 35, 19, 23] to generate synthetic datasets, to the best of our knowledge, we are the first ones to rely on a prefix tree to generate synthetic datasets.
Random sequence: We start at the root node that becomes the current node. Iteratively, we chose the next event from the current node with a probability weighted by the event frequencies (e.g., 40% of chance to pick “alphabloc” in the root node of Visualization of Electronic Health Record Sequences at Scale). This next event is appended to our random sequence. We iterate until the next event selected is “end of treatment.”At the end of this phase, we have a list of high-level events, and we also keep the list of the current nodes.
Adding Attributes: The nodes created at line 24 of LABEL:lst:aggregation by getHighLevelEventSequence contain a name, an age, a duration, and possible comorbidities. We generate the same attributes for our events at this stage. We generate the attributes backward, starting from the last event associated with the last node that we kept previously. Using the age distribution of the last node, we generate a random age that we set to the last created event. Same for the duration, and from the frequency of comorbidities, we generate a comorbidity pattern. Then, we proceed backward to the event list and tree nodes to create our events, updating the age according to the generated duration, propagating the comorbidity pattern (we do not change it yet), and picking a random duration again. This process is repeated until we reach the first event. Our high-level sequence is then complete.
Generating Low-Level Events: We first pick a plausible starting date at random for our low-level sequence. From the generated high-level sequence, we split events with duration into a sequence of atomic events (e.g., drug purchases) at regular intervals (15 days for each box of pills). Each low-level event is then given a date. We also remove the synthetic “interruption” and “no treatment” events, updating the date of the next concrete event accordingly. This phase generates a plausible sequence of events for one synthetic patient, similar to the CSV file Table 1.
@@
To be able to share synthetic data, according to national regulations, we should process the aggregated tree. The regulations are different for different countries. As fully anonymizing the aggregated tree is out of the scope of this article, we instead give readers interested in that topic a few references that discuss the potential threats of attackers, and the solutions to minimize them with the cost of lesser diversity in the aggregated data [24, 16, 8].
@@ Appendix C Additional Figures
Figure 12: Total computation time per size of the datasets compared to the number of threads; non progressive processing. The results follow the $\frac{1}{x}$ law; see also Figure 13.
Figure 13: Computational speed per size of the dataset compared to mono-threaded computations; non-progressive environment. The speed increases linearly with the number of CPU cores.
The two following figures are complementary to Figure 9 and Figure 10.
Figure 14: The 95% CIs, computed by Bootstrapping (BCA), of how fast the Hysteresis Sort stabilizes nodes compared to RSort by depth.
Figure 15: The 95% CIs, computed by Bootstrapping (BCA), of how fast the Hysteresis Sort stabilizes the tree compared to RSort by depth.
Figure 16: A heatmap showing how soon (in number of iterations) the Hysteresis Sort stabilizes nodes (categorized per their log-scaled frequency) compared to the RSort. Positive values along the y-axis mean that the Hysteresis Sort stabilizes nodes sooner than the RSort. The heatmap is computed for $\textit{threshold}=0$ and $\textit{chunkSize}=100K$, and considers only nodes whose depths are equal to or lower than 4, counting from 0. We use D3.js and its BuGn color scale.
Figure 17: The stability of the tree by depth and chunkSize values. We applied a small offset on the x-axis to distinguish overlapping points of two different chunkSize values. Each category has 33 MinSize values. The figure does not highlight clusters based on chunkSize, suggesting that this variable has no effect on how fast the tree stabilizes itself.
@@
ParcoursVis has been designed to visualize health records, but could be used for other kinds of event sequence data. We adapted it to visualize the MSNBC.com Anonymous Web Data, the largest public dataset mentioned in the related work [46]; it was retrieved from the UCI Machine Learning Repository [43]. It contains 989,925 sequences of website visits, with 18 types of web pages, i.e., event types, and the average number of visits per user is (average sequence length) 5.7.
@@
Figure 18: The MSNBC.com Anonymous Web Site dataset containing about 1 million web page visits visualized with ParcoursVis.

### v2

@@ 4.3.2. Results
Our evaluation shows that all nodes have a similar frequency from the first iteration to the last, regardless of MinSize and chunkSize, meaning that the frequency distribution of nodes converges fast for datasets like ours. As we shuffled the datasets, this result is expected for high values of chunkSize.
@@
We then look at how fast Hysteresis Sort and RSort stabilize in depth and relative position (9(a) and 9(b)). We normalized the iteration value to $[0,1]$ to compare the evaluations with different chunkSize. The distribution of the number of stabilized nodes and depths by iteration value is not normal. Thus, we cannot rely on parametric statistical tests. Following recommendations of (Dragicevic, 2016; Amrhein et al., 2019; Besançon and Dragicevic, 2019), we instead rely on estimation techniques with effect-size and confidence intervals (CIs). We bootstrap all the evaluation results per sorting strategy and per depth, resulting in $\textit{size}(\textit{MinSize})\times\textit{size}(\textit{chunkSize})$ different values per bootstrap. Depth starts from 0, which corresponds to the children of the root, as the root is, in our case, an abstract object that we do not visualize in ParcoursVis.
@@
9(a) shows that the prefix tree is stable after the progressive algorithm processed $80\%$ (at the right edge of the 95% CI) of the dataset at the maximum tested $\textit{depth}=4$ using the Hysteresis Sort, compared to approximately 90% using RSort. Our results (9(b)) show statistically significant differences that Hysteresis Sort and RSort start to diverge at depth=1, with a large effect size for $\textit{depth}\in\{1,2\}$. Because deeper depths depend on shallower ones, the results of RSort for $\textit{depth}=1$ strongly impact the deeper results. This seems less the case for Hysteresis Sort (9(a)).
@@
We then looked at the stability per node, categorized by their depths. 9(b) shows strong evidence that Hysteresis Sort outperforms RSort at stabilizing nodes sooner for all depths. In addition, we looked at the stability of nodes categorized by their frequencies. We computed, per node, the difference in stability of Hysteresis Sort compared to RSort for the evaluation, set with $\textit{MinSize}=0$ and $\textit{chunkSize}=100\textrm{k}$, which are the minimum values we tested. We see that Hysteresis Sort and RSort behave the same for most nodes, which we expected, as most nodes do not have siblings with similar frequencies (see $\textit{iteration}=0$). Second, the sorting strategy mostly impacts nodes with low frequency (lower than 200, i.e., representing less than 1% of our population). Third, while Hysteresis Sort strongly improves the stability of most of the remaining nodes, it can degrade the stability of some small nodes for a few iterations. Overall, we conclude on the effectiveness of Hysteresis Sort compared to RSort. With our datasets and standard setting of MinSize=50, the tree is completely stable when 40% of the dataset has been processed (9(c)).
@@
Finally, we wanted to see if the variables MinSize and chunkSize have an effect on the stability. 9(c) suggests that MinSize strongly impacts the stability of the positions of the nodes by removing small changes. Because the dataset is shuffled, large nodes tend to stabilize sooner than smaller ones, explaining the results. However, this figure shows that MinSize has a stronger impact on the Hysteresis Sort than on the RSort. We do not find any effect of the chunkSize variable on the stability of nodes.
4.4. User Feedback
@@
We collected user feedback from both use cases introduced in Section 3.2, enabling participants to visualize datasets with an unprecedented level of scalability with ParcoursVis. We provide usability insights, focusing on its progressive aspect when used by experts. Participants were able to perform tasks comparable to those in EventFlow systems with no perceived latency. This study has been approved by our organization’s Ethical Committee.
@@
Experts were excited to be able to explore such large datasets to eventually find unexpected cases. We interviewed a referring doctor to explore the non-cancerous prostate adenoma dataset and received positive feedback. He reported that his wife thought he had a new video game because he was spending hours exploring the data interactively. He also collected many insights and surprising findings that need inquiry now. For example, he thought that surgery would cure the patients when, in reality, many patients ($18\%$) need to resume treatment and sometimes undergo another surgery. He also formulated hypotheses that require validation through further examination of patients’ cases (e.g., long drug treatments on severe cases may weaken the urinary tract).
We interviewed six additional healthcare professionals on the ED dataset: four emergency physicians (two with extensive experience in ED operations and two focused on improving organizational efficiency), and two clinical research data analysts. We also interviewed a researcher in computer science focused on high-fidelity simulation of patient flow in emergency departments. During the study, we used data from three hospitals for a total of 476,109 visitors for reasons of data quality and comparability. We asked them to complete some of the tasks listed in Section 3.2 using a think-aloud approach, followed by a semi-structured interview; all sessions were audio-recorded.
They all succeeded in performing the tasks. None of our participants mentioned any latency issues, not even performance issues. They were able to focus on their professional questions.
@@
Additionally, we used ParcoursVis to answer several data-related questions during project meetings. For example, we were able to quickly characterize and select important pathways to optimize, such as the most frequent ones, the most time-consuming ones, and the most “undesirable” ones, e.g., revisiting the ICU, or the emergency service in a short time.
5. Discussion and Future Work
5.1. Performance Analysis
@@
The analysis described in Section 4.2 validates the linear complexity of our aggregation algorithm, both in the number of threads and low-level events. We aggregate 10 million synthetic patients ($\approx 180\textrm{M}$ low-level events) on our dedicated computer in less than one second. ParcoursVis is meant to be used on a server shared by multiple users; aggregating sequences in the wild takes longer. Moreover, our evaluation relies on short sequences on which we applied a small number of filters. Longer sequences and more filters cause slowdowns. PVA solves these issues with almost no overhead.
@@
We filter small nodes to avoid confetti visualizations, but users can configure it. In our evaluation using $\textit{MinSize}=50$ (i.e., 0.0025% of the two million patients dataset), the view aggregating the dataset stabilizes at around 40% of the computation time for the deepest nodes using Hysteresis Sort, 25% sooner than RSort (9(c)). The MinSize variable thus can strongly increase the speed at which users can work on their data if they are not interested in small nodes.
5.2. Stability
@@
We evaluated the stability of aggregated trees for one use case with an inertia of $20/1080$. For that configuration, we show Hysteresis Sort is better than RSort. We believe this generalizes to other datasets, although the stabilization times may vary depending on the distribution of events. Our Hysteresis Sort algorithm relies on a modified version of the bubble sort algorithm with a theoretical worst-case quadratic complexity. In practice, our sort time is linear because we use a limited number of types to be able to assign them to color effectively (as explained in Section 2.2.2). The scalability of ParcoursVis and ESS, in general, remains limited by the number of event types they can handle (Wang and Laramee, 2022).
5.3. Evaluation Criteria for Progressive Visualization
@@
Progressive systems are relatively new, and we need more evaluation criteria to assess their usability (Richer et al., 2024a). We did not consider comparing the usability of our system in its current state to a non-progressive system, given the latency we measured, which makes non-progressive systems completely unusable (measures are in the Supplementary Material). During our expert interviews, participants were unaware that the tool used progressive algorithms and did not notice it; we consider this lack of mention a usability success for PVA. Yet, it is a new qualitative evaluation criterion.
@@
As explained in Section 2.3, we rely on stability as a proxy for convergence. For update durations limited to tens of seconds, we believe instability is an acceptable proxy for uncertainty, and our expert users did not express any confusion or trust issues about it. Longer update times may need explicit uncertainty visualization, though.
5.4. Improving the Scalability of EHR Visualizations
@@
5.5. Future Work
Each new medical study requires new features, and we are planning to extend ParcoursVis to make it more generic and applicable to new use cases with little programming, providing a larger set of aggregation options as a library.
We also plan to support datasets updated continuously. That would require certain adjustments to the system, specifically running regular pre-processing of the updated database. ParcoursVis should be able to restart immediately when a new dataset is available, but it should also avoid interrupting ongoing explorations.
ParcoursVis is currently being applied to a new pathology, prostate cancer, using data from the SNDS (SNDS, 2023). While we currently handle only a limited number of attributes per patient, the dataset provides thousands of attributes. To enable richer exploration, we aim to develop methods to manage these attributes dynamically and on demand, including the ability to display their distributions, scaling ParcoursVis to effectively support another effort: a large number of attributes and types. One way to achieve this, building on the idea introduced in Section 3.4.2, is to keep the patient identifiers in the aggregation tree and compute patient-level aggregations on the fly from the node patient list. We would rely on an external database such as DuckDB (Raasveldt and Mühleisen, 2019) to compute the distributions or aggregated values. This approach would scale ParcoursVis to handle not only a large number of patients but also a large variety of attributes and types, without overloading bandwidth or memory.
6. Conclusion
We presented ParcoursVis, a Progressive Visual Analytics tool we designed to explore patients’ care pathways at scale. We described its features, including the history navigation and comparison, which are novel for EHR visualization systems. We described its progressive architecture, scalable aggregation algorithm, and its stabilizing sorting algorithm. Using Richer et al. (Richer et al., 2024b) scalability model, we applied quantitative evaluations showing the aggregation algorithm scales linearly with the number of events and inversely with the number of threads, always meeting interactive latency limits. We also showed that our Hysteresis sort improves the stability of the Icicle tree visualization. ParcoursVis can be used to explore patients’ treatments and conditions aggregated at the scale of the largest countries to improve public health based on data.
For a dataset similar to our benchmark scaled to China’s population, non-progressive aggregation on our machine would take 27 s to complete. In contrast, our PVA architecture ensures updates in under 2 s, with most tree nodes stabilizing within 10 s, enabling exploration of datasets at the scale of the largest countries.
Adapting visualizations to become progressive and scalable still needs experience and novel software architectures; this article explained how we addressed important issues. Increasing the scalability of aggregated sequence visualization by three to five orders of magnitude, depending on the measure, allows addressing new sets of problems from larger medical data. By providing our system as open-source, we want to push this research field to a wider audience, relying on PVA to provide scalability and improve public health worldwide.
@@ 9 Histogram age_distribution; // age distribution
Listing 1: Structure of a tree Node
The tree contains nodes (1), each node $N_{i}$ representing a pathway $P_{i}$ from the root to $N_{i}$:
@@
  • children: The list of events following $P_{i}$. For each event $E_{j}$, we export its frequency, i.e.probability of $P({E_{j}}|{P_{i}})$.
@@
From this tree, we can generate a new patient in three phases: (1) generating a random sequence of high-level events from the tree, (2) adding realistic attributes to the high-level events, and (3) generating a plausible low-level event sequence from the high-level events. While it is common to rely on statistics and machine learning models (Goncalves et al., 2020; Walonoski et al., 2020; Ping et al., 2017; Guan et al., 2021; Kokosi and Harron, 2022) to generate synthetic datasets, to the best of our knowledge, we are the first ones to rely on a prefix tree to generate synthetic datasets.
Random sequence: We start at the root node that becomes the current node. Iteratively, we chose the next event from the current node with a probability weighted by the event frequencies (e.g.40% of chance to pick “alphabloc” in the root node of Figure 1). This next event is appended to our random sequence. We iterate until the next event selected is “end of treatment.”At the end of this phase, we have a list of high-level events, and we also keep the list of the current nodes.
Adding Attributes: The nodes created at line 24 of Listing 2 by getHighLevelEventSequence contain a name, an age, a duration, and possible comorbidities. We generate the same attributes for our events at this stage. We generate the attributes backward, starting from the last event associated with the last node that we kept previously. Using the age distribution of the last node, we generate a random age that we set to the last created event. Same for the duration, and from the frequency of comorbidities, we generate a comorbidity pattern. Then, we proceed backward to the event list and tree nodes to create our events, updating the age according to the generated duration, propagating the comorbidity pattern (we do not change it yet), and picking a random duration again. This process is repeated until we reach the first event. Our high-level sequence is then complete.
Generating Low-Level Events: We first pick a plausible starting date at random for our low-level sequence. From the generated high-level sequence, we split events with duration into a sequence of atomic events (e.g.drug purchases) at regular intervals (15 days for each box of pills). Each low-level event is then given a date. We also remove the synthetic “interruption” and “no treatment” events, updating the date of the next concrete event accordingly. This phase generates a plausible sequence of events for one synthetic patient, similar to the CSV file Table 1.
@@
To be able to share synthetic data, according to national regulations, we should process the aggregated tree. The regulations are different for different countries. As fully anonymizing the aggregated tree is out of the scope of this article, we instead give readers interested in that topic a few references that discuss the potential threats of attackers, and the solutions to minimize them with the cost of lesser diversity in the aggregated data (Li et al., 2007; Fung et al., 2010; Brickell and Shmatikov, 2008).
@@ Appendix C Additional Figures
Figure 10. Total computation time per size of the datasets compared to the number of threads; non progressive processing. The results follow the $\frac{1}{x}$ law; see also Figure 11.
Figure 11. Computational speed per size of the dataset compared to mono-threaded computations; non-progressive environment. The speed increases linearly with the number of CPU cores.
The two following figures are complementary to 9(a) and 9(b).
Figure 12. The 95% CIs, computed by Bootstrapping (BCA), of how fast the Hysteresis Sort stabilizes nodes compared to RSort by depth.
Figure 13. The 95% CIs, computed by Bootstrapping (BCA), of how fast the Hysteresis Sort stabilizes the tree compared to RSort by depth.
Figure 14. A heatmap showing how soon (in number of iterations) the Hysteresis Sort stabilizes nodes (categorized per their log-scaled frequency) compared to the RSort. Positive values along the y-axis mean that the Hysteresis Sort stabilizes nodes sooner than the RSort. The heatmap is computed for $\textit{threshold}=0$ and $\textit{chunkSize}=100K$, and considers only nodes whose depths are equal to or lower than 4, counting from 0. We use D3.js and its BuGn color scale.
Figure 15. The stability of the tree by depth and chunkSize values. We applied a small offset on the x-axis to distinguish overlapping points of two different chunkSize values. Each category has 33 MinSize values. The figure does not highlight clusters based on chunkSize, suggesting that this variable has no effect on how fast the tree stabilizes itself.
@@
ParcoursVis has been designed to visualize health records, but could be used for other kinds of event sequence data. We adapted it to visualize the MSNBC.com Anonymous Web Data, the largest public dataset mentioned in the related work (Vrotsou and Nordman, 2019); it was retrieved from the UCI Machine Learning Repository (UCI ML, [n. d.]). It contains 989,925 sequences of website visits, with 18 types of web pages, i.e., event types, and the average number of visits per user is (average sequence length) 5.7.
@@
Figure 16. The MSNBC.com Anonymous Web Site dataset containing about 1 million web page visits visualized with ParcoursVis.